# A Simple and Explainable Model for Park-and-Ride Car Park Occupancy Prediction

### v1

@@ Abstract
The proposed model successfully identifies the regular, periodic nature of commuter parking behaviour, where vehicles arrive in the morning and depart in the afternoon. It operates using aggregate data, eliminating the need for individual tracking of arrivals and departures. The model’s predictive and now-casting capabilities are demonstrated through real-world data from car parks in the Barcelona Metropolitan Area. A simple model extension furthermore enables the prediction of when a car park will reach its occupancy limit and estimates the additional spaces required to accommodate such excess demand. Thus, beyond forecasting, the model serves as a valuable tool for evaluating interventions, such as expanding parking capacity, to optimize park-and-ride facilities.
@@
Car Park, Occupancy Prediction, Demand estimation, Park-and-ride, Commuter Behaviour
@@ 1 Introduction
In this work, we use data collected from a set of eight park-and-ride facilities located in the vicinity of Barcelona, Catalonia. These car parks primarily serve commuters on working days, who typically park their vehicles in the morning before transferring to public transport for work or study in the city. In the afternoon, they return to retrieve their cars and drive home.
@@
To achieve these goals, we will model car arrivals and departures separately, each using a different truncated normal distribution. This approach allows us to use aggregate data without the need to monitor individual arrivals and departures. Before presenting the proposed model in Section 4, we first review relevant literature in Section 2. In Section 3, we describe the dataset used in our study. Section 5 then analyzes the model’s ability to fit the observed data and evaluates its performance in both prediction and nowcasting tasks. Finally, we conclude with a discussion of our findings in Section 6.
@@ 2 Related work
The literature about smart parking solutions is vast and includes different disciplines and types of problems as can be observed in this survey paper [1]. Here, we focus on a specific problem in this area: Modeling car park occupancy. This has been treated in the literature mostly only as a prediction problem (see [2] for a recent review on the subject). Two different types of prediction problems can be distinguished: on-street and off-street (e.g. in garages and parking lots) parking prediction. We are aiming at the latter.
@@
Notable works for car park occupancy prediction include [3], [4], and [5]. Reference [3] compares the performance of regression trees, support vector regression and a neural network with a single hidden layer for sensor-enabled car parks from San Francisco (US) and Melbourne (Australia). The study found that the least computationally intensive algorithm of the three, the regression trees with the history of the occupancy rates and the weekday as input features performed best.
@@
Recurrent neural networks with more hidden layers have then been employed by [4] to predict car park occupancy rates of 29 car parks in Birmingham, UK.
Furthermore, the authors of [5] also give an overview of different parking occupancy prediction methods. Their study compares methods based on linear regression, support vector machine (SVM), neural networks and auto-regressive integrated moving average (ARIMA) which are applied to data from four car parks in three Chinese cities. The effect of distinguishing between weekend and weekday patterns is also analyzed. The study concludes that SVM offers stable and accurate prediction performance for almost all types and scales of parking lots while the distinction between weekday and weekend only led to improvements for medium size parking lots and office space parking.
In the context of commuter behaviour, [6] analyzed the relationship between parking space availability and commuters’ departure time choices using survey data. This study also includes a model with a bell-shaped distribution for the perceived value of choosing a specific departure time, which produces departure distributions similar to the ones we use in our models.
Patterns with multiple daily peaks, more complex than the ones we observe here have been analyzed in [7] with the aim of predicting parking space availability in a more generic setting without using an underlying behavioural model. The same aim has been pursued in [8] which models individual parking lot occupancy as a two-state stochastic process. Similar patterns have been analysed recently in [9] where the number of parked car-sharing vehicles in specific areas in the Milan metropolitan area are predicted with a Maximum Entropy modelling framework which allows identifying extreme events such as bad weather conditions or strikes.
Here, we take a slightly more ambitious approach by aiming to understand the underlying behavioural patterns. In particular, we compare our method with simple prediction techniques based on average activity profiles, as used in [10] and [11]. Previous studies have shown that historical data can effectively predict future parking occupancy rates, even with limited data [12].
This study (as does the above-mentioned reference [5]) uses among outer algorithms linear regression as prediction algorithm. We will use this method as an additional baseline to compare the predictive power of our modelling approach.
For cases where (contrary to our case) individual arrival and departure times are available a related modelling approach based on queuing models with non-homogeneous arrival rates was undertaken in [13]. Using data from truck parking lots the authors found that a queuing model of parking dynamics and a model-based prediction method can provide real-time probabilistic forecasts of future parking occupancy.
Behavioural decision models for park-and-ride facilities utilization have been analyzed in [14], whose authors use survey data to analyze how personal attributes, travel characteristics, and user intentions influence the decisions to choose or not a Park-and-Ride trip. Shortage of parking space is found to be among the top three reasons not to use Park-and-Ride facilities. This shows the importance of not only correctly predicting the demand for parking space but also the missing amount of parking space which cannot be fulfilled by the current infrastructure. None of the above-mentioned studies has tackled this problem, which our behavioural model can address.
@@ 3 Dataset Description
We use data provided by the metropolitan transport authority  [^1] Autoritat de Transport Metropolità (ATM): https://www.atm.cat/en/atm. of Barcelona (Catalonia) which has installed sensors in different park-and-ride parking locations with the aim to better understand their usage patterns and to develop a long-term strategy that allows to improve the interurban mobility from and to Barcelona. Each of the car parks is located in the vicinity of a regional train station which allows people to park their cars and change directly there to the public transportation system, working as a bridge between both public and private transportation means [15].
Figure 1: Locations of the park-and-ride car parks analyzed, all of them in the vicinity of Barcelona (Catalonia). Two of the locations, Martorell and San Quirze, have not been used due to errors in the data collection process.
Data from a total amount of 10 parking lots have been provided. Their geographical location is depicted in Figure 1. Data is continuously gathered from physical sensors placed in the entrances and exits of the car parks and stored in occupancy counters. An entrance of a vehicle increases the occupancy of the particular parking, and a departure of a vehicle decreases that counter. No data about the precise physical location where a vehicle park is recorded. This occupancy data is sampled in 30-minute internals from every parking during a 3-month period from January 1st 2020 0:00 to April 1st 2020 0:00. Note that Catalonia entered a COVID-19-induced nationwide lockdown on March 15, 2020, and the car park usage decreased significantly. We will omit the data posterior to this date in the subsequent analysis. We furthermore have removed data from two car parks (Martorell and San Quirze) due to defective sensors and storage errors which led to irregular readings.
Figure 2 gives an example of the raw data from one of the car parks (Vilanova). The observed occupancy daily pattern is the aggregate result of arrivals and departures to the car park location. In the subsequent analysis, we have manually removed days where the occupancy profiles show errors due to sensor or storage failures. An example of such days are the days from February 7 to February 9 in Figure 2 or days with are influenced by holidays or vacation periods (e.g days 1 to 3 and 6 of January in Figure 2, both periods together with the also removed COVID-19 lockdown period are shown with a grey background.
@@
In what follows we aggregate the data in three groups: weekdays from Monday to Thursday, Fridays, and weekends (Saturdays and Sundays)
@@
Finally in Figure 4 we look at how close the different car parks get to their capacity limit (indicated by the red vertical line). For every day (excluding the days we have filtered out as mentioned above) we record the maximum occupancy of the car park and depict the corresponding distributions. Four of the car parks (Sant Sadurni, Sant Boi, Quatre Camins and Mollet) reach their maximal occupancy limit regularly during weekdays (orange bars) and also Fridays (blue bars). In the case of the example we used above, the Vilanova car park, this is not the case. We will use this distinction and develop a model extension for the stations where the capacity limit is reached.
Figure 4: Histogram of the observed maximal occupancies for the days between 01-01-2020 and 01-04-2020 aggregated by weekdays (orange), Fridays (blue) and weekends (green) for the eight car parks used in this study. Red lines indicate the maximal capacities of the car parks.
4 Modeling Car Park Occupancy
@@
We introduce two types of models: a basic model which assumes an unlimited amount of free parking spaces, and an extended model with occupancy limit. The former is suitable for parking lots which do not fill up completely whereas the latter takes into account that a parking lot may reach its capacity limit and cars arriving after this moment will have to find parking space elsewhere.
@@ 4.1 Basic Model (Unlimited Parking Spaces)
The proposed model accounts for two event types: car arrivals and departures. For simplicity, these events are assumed to be independent. Each process is modelled as a random variable following a (truncated) normal distribution. The truncation is necessary to constrain the arrival and departure times to occur within the period corresponding to one day (which we normalize to the interval $[0,1]$). The probability density functions for arrivals $\phi_{a}$ and departures $\phi_{d}$ are thus:
@@
where $\phi(\cdot)$ is the probability density function of the standard normal distribution and ${\displaystyle\Phi(\cdot)}$ its cumulative distribution function (CDF). We denote the truncated CDFs as $\Phi_{a}(t;\mu_{a},\sigma_{a})$ for the arrival and $\Phi_{d}(t;\mu_{d},\sigma_{d})$ for the departure process, which are determined by parameters $(\mu_{a},\sigma_{a})$ and $(\mu_{d},\sigma_{d})$ respectively.
@@
where $\mathbbm{1}_{\{\cdot\}}$ is the indicator function that takes value 1 when the condition is true and 0 otherwise. To know the occupancy of a car park at a given time $t$ we calculate the difference between all arrivals and departures happened before.
@@
Figure 5: Example of the TN model with data from the Vilanova car park. (top) probability density function (PDF) of car arrival (blue) and departure times (red), (bottom) corresponding CDFs (dashed lines) and car park occupancy curve (red think solid line).
@@ 4.2 Model with Occupancy Limit
We can extend the basic TN model to account for a capacity limit in the parking lot by incorporating a new parameter into the model. Consider the total number of cars arriving in one day and let $\tau$ be the fraction of those cars that fit in the parking lot. We define the Truncated Normal with Limit (TNL) model as
@@
with $\text{with }~{}\bar{\Phi}_{a}(t;\mu_{a},\sigma_{a},\tau)=\min\left(\frac{\Phi_{a}(t;\mu_{a},\sigma_{a})}{\tau},1\right)$ (5)
@@
The TNL model allows, not only to characterize the moment $t_{L}$ when the capacity limit is reached (i.e. $\Phi_{a}(t_{L};\mu_{a},\sigma_{a})=\tau$), but also to have an estimate of the number of cars which did not find a free parking space. This is very relevant in scenarios where informed decision-making is required to determine policies to resize crowded parking lots.
Figure 5 illustrates the TNL model for a day when the parking lot capacity limit was reached between 8 and 8:30 (vertical dashed line in cyan). The top panel shows an example of a density function for arrivals (in blue) and departures (in red). The arrival distribution is truncated when the parking fills up. The dashed-dotted black line of the arrival distribution corresponds to the estimated density of cars arriving afterwards which no longer fit in the parking lot. The bottom panel shows the corresponding modified CDFs $\bar{\Phi}_{a}$ (dashed line in blue) and $\Phi_{d}$ (dashed line in red) together with the difference $\bar{\Phi}_{a}-\Phi_{d}$ (red solid line), as the aggregate result of cars parked in the lot. Again, the proportion of surplus cars is indicated by the dashed-dotted line.
@@
4.3 Optimizing Model Parameters from Observed Data

### v2

@@ Abstract
The proposed model successfully identifies the regular, periodic nature of commuter parking behaviour, where vehicles arrive in the morning and depart in the afternoon. It operates using aggregate data, eliminating the need for individual tracking of arrivals and departures. The model’s predictive and nowcasting capabilities are demonstrated through real-world data from car parks in the Barcelona Metropolitan Area. A simple model extension furthermore enables the prediction of when a car park will reach its occupancy limit and estimates the additional spaces required to accommodate such excess demand. Thus, beyond forecasting, the model serves as a valuable tool for evaluating interventions, such as expanding parking capacity, to optimise park-and-ride facilities.
@@
Car Park, Parking Lot, Occupancy Prediction, Demand estimation, Park-and-ride, Commuter Behaviour
@@ 1 Introduction
In this work, we use data collected from a set of eight park-and-ride facilities located in the vicinity of Barcelona, Catalonia. These car parks [^1] Note: throughout this paper, we use the term “car park” following British English usage, which corresponds to “parking lot” in American English. primarily serve commuters on working days, who typically park their vehicles in the morning before transferring to public transport for work or study in the city. In the afternoon, they return to retrieve their cars and drive home.
@@
To achieve these goals, we will model car arrivals and departures separately, each using a different truncated normal distribution. This approach allows us to use aggregate data without the need to monitor individual arrivals and departures. Before presenting the proposed model in Section 4, we first review relevant literature in Section 2. In Section 3, we describe the dataset used in our study. Section 5 then analyses the model’s ability to fit the observed data and evaluates its performance in both prediction and nowcasting tasks. Finally, we conclude with policy recommendations in Section 6 and a discussion of our findings in Section 7.
@@ 2 Related work
The literature on smart parking solutions is extensive, spanning multiple disciplines and problem types, as highlighted in the survey by Lin et al. [1]. Our work focuses on a specific aspect within this area: modelling car park occupancy. Traditionally, this has been approached primarily as a prediction task (see [2] for a recent overview). Prediction problems can be classified into two main categories: on-street parking and off-street parking (e.g., garages and dedicated parking lots). Our research specifically addresses the latter.
This section is organised around three main limitations identified in existing studies: limited model interpretability, challenges in handling aggregate data instead of individual vehicle arrivals and departures, and insufficient capacity to account for unmet demand, i.e., reasoning about counterfactual scenarios.
@@
Notable studies that focus specifically on car park occupancy prediction include [3], [4], [5], and [6]. In [3], the authors compared regression trees, support vector regression, and single-layer neural networks for predicting occupancy in sensor-equipped on-street parking in San Francisco (US) and Melbourne (Australia). They concluded that regression trees, using historical occupancy rates and weekdays as input features, provided the best performance with the least computational cost.
@@
Building on this, recurrent neural networks with multiple hidden layers were employed by [4] to predict occupancy rates in 29 off-street car parks located in Birmingham, UK.
In another comparative analysis, Zhao et al. [5] reviewed various prediction methods, including linear regression, support vector machines (SVM), neural networks, and auto-regressive integrated moving average (ARIMA). Their evaluation, based on data from four car parks across three Chinese cities, also explored the impact of differentiating between weekdays and weekends. They found that SVM consistently provided stable and accurate predictions across nearly all car park types and sizes, while distinguishing weekdays from weekends improved results mainly for medium-sized and office-related parking lots.
Similarly, Awan et al. [6] analysed sensor-collected occupancy data from Santander, a smart city in Spain, over a period of nine months. Their study concluded that simpler algorithms, such as random forests, systematically outperformed more complex deep learning models, such as neural networks.
All these approaches share the three limitations we aim to address in our work.
Several other works relax the assumptions regarding data requirements. However, these approaches still face important limitations in terms of interpretability and their ability to model unmet demand.
Chen [7] consider different levels of data aggregation to predict parking space availability without relying on an explicit behavioural model. They analyse complex usage patterns, including multiple daily peaks, which are more intricate than those addressed in our work.
Ji et al. [8] propose a wavelet neural network (WNN) model for short-term parking prediction, demonstrating higher accuracy and computational efficiency compared to traditional methods using real-world data from Newcastle, UK.
Chawathe [9] introduce a data-driven predictive approach that leverages historical occupancy rates to estimate current and future availability. A key advantage of this method is that it requires minimal or no additional instrumentation of parking facilities, while still achieving sufficient accuracy for practical guidance.
Fokker et al. [10] conduct long-term (6-months ahead) off-street occupancy forecasting for 57 parking facilities in Amsterdam. Their models incorporate external factors such as weather, local events, parking tariffs, and public transport changes. While this work incorporates rich contextual data, the models used offer limited interpretability.
Vakrinou et al. [11] address both short- and long-term occupancy forecasting across multiple off-street parking facilities in central Athens, using a multisource dataset that includes traffic, public transport operations, and weather conditions. Their approach employs LSTMs, Transformers, and XGBoost. Although XGBoost provides a degree of interpretability, the models are not grounded in an explicit behavioural framework, and cannot be used to reason about unmet demand.
Table 1: Overview of limitations in existing work on car park occupancy modelling. The table summarises whether each group of studies provides interpretable models, operates without requiring high-granularity data (e.g., individual arrival and departure times), and whether they can account for unmet demand. Our proposed framework addresses all three aspects.
Related work Provides interpretable models Does not require high-granularity data Able to handle unmet demand
Zheng et al. [3], Camero et al. [4], Zhao et al. [5], Awan et al. [6]
Chen [7], Ji et al. [8], Chawathe [9], Zhao et al. [5], Fokker et al. [10], Vakrinou et al. [11]
Tavafoghi et al. [12], Daniotti et al. [13], Schneble and Kauermann [14]
Some recent works introduce interpretable models, but they still rely on fine-grained data and are unable to reason about unmet demand.
Tavafoghi et al. [12] propose a queuing framework with non-homogeneous arrival rates for predicting occupancy in truck parking lots. Their method leverages individual arrival and departure times to estimate the service time distribution, assuming four non-overlapping heterogeneous populations of arrivals, each with a time-independent distribution. They also assume that no arriving vehicle finds the parking lot full, i.e., there are infinitely many parking spots.
Schneble and Kauermann [14] model on-street individual parking lot occupancy in the city centre of Melbourne as a two-state stochastic process using a semi-Markov framework. While the model is interpretable and statistically grounded, it requires detailed sensor-level data at the level of individual parking spots.
Similarly, Daniotti et al. [13] analyses patterns in the number of parked car-sharing vehicles across specific areas in the Milan metropolitan region. They employ a Maximum Entropy modelling approach that enables the identification of rare or extreme events, such as weather anomalies or transit strikes. Despite its interpretability, the method also depends on high-resolution, disaggregated data.
Table 1 summarises the main limitations of the reviewed approaches along the three dimensions that motivate our work: model interpretability, ability to operate with aggregate data, and capacity to reason about unmet demand.
Our proposed framework is inspired by prediction techniques based on average activity profiles, such as those used in [15, 16]. Although these methods were originally developed for non-parking mobility data, they illustrate how temporal regularities in aggregate behaviour can be leveraged for forecasting. Similar ideas have been applied in the parking domain, where historical data has proven effective in predicting future occupancy levels even with limited data availability [9].
We end this section by reviewing conceptually related work which considers behavioural aspects of parking demand using questionnaire data. For example, Xue et al. [17] models commuter departure time choices under parking capacity constraints using bounded rationality principles. While their goal is not occupancy prediction per se, their approach highlights the role of limited parking availability in shaping temporal demand patterns and motivates the use of bell-shaped temporal distributions, a key component of our modelling framework.
Similarly, Zhao et al. [18] analyses parking behaviour in the context of Park-and-Ride facilities, using survey data to study how personal attributes, travel characteristics, and user intentions influence the decision to use such infrastructure. Their findings reveal that shortage of parking space is among the top three reasons users avoid Park-and-Ride, emphasising the importance of modelling not only observed demand but also unmet demand.
@@ 3 Dataset Description
We use data provided by the metropolitan transport authority  [^2] Autoritat del Transport Metropolità (ATM): https://www.atm.cat/en/atm. of Barcelona (Catalonia) which has installed sensors in different park-and-ride parking locations with the aim to better understand their usage patterns and to develop a long-term strategy that allows to improve the interurban mobility from and to Barcelona. Each of the car parks is located in the vicinity of a regional train station, which allows people to park their cars and change directly there to the public transportation system, working as a bridge between both public and private transportation means [19].
Figure 1: Locations of the park-and-ride car parks analysed, all of them in the vicinity of Barcelona (Catalonia). Two of the locations, Martorell and San Quirze, have not been used due to errors in the data collection process. Scale indicates a 5km distance.
Data from a total of 10 parking lots have been provided. Their geographical location is depicted in Figure 1. Data is continuously gathered from physical sensors placed in the entrances and exits of the car parks and stored in occupancy counters. The entrance of a vehicle increases the occupancy of the particular car park, and the departure of a vehicle decreases that counter. No data about the precise physical location where a vehicle is parked is recorded. This occupancy data is sampled in 30-minute intervals from every parking during a 3-month period from January 1st 2020, 0:00 to April 1st 2020, 0:00. Note that Catalonia entered a COVID-19-induced nationwide lockdown on March 15, 2020, and the car park usage decreased significantly. We will omit the data posterior to this date in the subsequent analysis. We have furthermore removed data from two car parks (Martorell and San Quirze) due to defective sensors and storage errors, which led to irregular readings.
Figure 2 gives an example of the raw data from one of the car parks (Vilanova). The observed occupancy daily pattern is the aggregate result of arrivals and departures to the car park location. In the subsequent analysis, we have manually removed days where the occupancy profiles show errors due to sensor or storage failures. An example of such days are the days from February 7 to February 9 in Figure 2 or days that are influenced by holidays or vacation periods (e.g days 1 to 3 and 6 of January in Figure 2, both periods together with the also removed COVID-19 lockdown period are shown with a grey background.
@@
In what follows, we aggregate the data in three groups: weekdays from Monday to Thursday, Fridays, and weekends (Saturdays and Sundays)
@@
Finally, in Figure 4 we look at how close the different car parks get to their capacity limit (indicated by the red vertical line). For every day (excluding the days we have filtered out as mentioned above), we record the maximum occupancy of the car park and depict the corresponding distributions. Four of the car parks (Sant Sadurni, Sant Boi, Quatre Camins and Mollet) reach their maximal occupancy limit regularly during weekdays (orange bars) and also Fridays (blue bars). In the case of the example we used above, the Vilanova car park, this is not the case. We will use this distinction and develop a model extension for the stations where the capacity limit is reached.
Figure 4: Histogram of the observed maximal occupancies for the days between 01-01-2020 and 01-04-2020 aggregated by weekdays (orange), Fridays (blue) and weekends (green) for the eight car parks used in this study (bin sizes =9=9 in all panels). Vertical red lines at the right of the panels indicate the maximal capacities of the car parks. The maximum capacity was reached in SantSadurni (19 times), SantBoi (39 times) and in QuatreCamins (45 times).
4 Modelling Car Park Occupancy
@@
We introduce two types of models: a basic model which assumes an unlimited amount of free parking spaces, and an extended model with occupancy limit. The former is suitable for parking lots which do not fill up completely, whereas the latter takes into account that a parking lot may reach its capacity limit, and cars arriving after this moment will have to find parking space elsewhere.
@@ 4.1 Basic Model (Unlimited Parking Spaces)
The proposed model accounts for two event types: car arrivals and departures. For simplicity, these events are assumed to be independent. Each process is modelled as a random variable following a (truncated) normal distribution. The truncation is necessary to constrain the arrival and departure times to occur within the period corresponding to one day (which we normalise to the interval $[0,1]$). The probability density functions for arrivals $\phi_{a}$ and departures $\phi_{d}$ are thus:
@@
where $\phi(\cdot)$ is the probability density function of the standard normal distribution and $\Phi(\cdot)$} its cumulative distribution function (CDF). We denote the truncated CDFs as $\Phi_{a}(t;\mu_{a},\sigma_{a})$ for the arrival and $\Phi_{d}(t;\mu_{d},\sigma_{d})$ for the departure process, which are determined by parameters $(\mu_{a},\sigma_{a})$ and $(\mu_{d},\sigma_{d})$ respectively.
@@
where $\mathbbm{1}_{\{\cdot\}}$ is the indicator function that takes the value 1 when the condition is true and 0 otherwise. To know the occupancy of a car park at a given time $t$, we calculate the difference between all arrivals and departures that happened before.
@@
Figure 5: Example of the TN model with data from the Vilanova car park. (top) probability density function (PDF) of car arrival (blue) and departure times (red), (bottom) corresponding CDFs (dashed lines) and car park occupancy curve (red thick solid line).
@@ 4.2 Model with Occupancy Limit
We can extend the basic TN model to account for a capacity limit in the parking lot by incorporating a new parameter into the model. Consider the total number of cars arriving in one day, and let $\tau$ be the fraction of those cars that fit in the parking lot. We define the Truncated Normal with Limit (TNL) model as
@@
$\text{with }~\bar{\Phi}_{a}(t;\mu_{a},\sigma_{a},\tau)=\min\left(\frac{\Phi_{a}(t;\mu_{a},\sigma_{a})}{\tau},1\right)$ (5)
@@
The TNL model allows not only to characterise the moment $t_{L}$ when the capacity limit is reached (i.e. $\Phi_{a}(t_{L};\mu_{a},\sigma_{a})=\tau$), but also to have an estimate of the number of cars which did not find a free parking space. This is very relevant in scenarios where informed decision-making is required to determine policies to resize crowded parking lots.
Figure 6 illustrates the TNL model for a day when the parking lot capacity limit was reached between 8 and 8:30 (vertical dashed line in cyan). The top panel shows an example of a density function for arrivals (in blue) and departures (in red). The arrival distribution is truncated when the parking fills up. The dashed-dotted black line of the arrival distribution corresponds to the estimated density of cars arriving afterwards, which no longer fit in the parking lot. The bottom panel shows the corresponding modified CDFs $\bar{\Phi}_{a}$ (dashed line in blue) and $\Phi_{d}$ (dashed line in red) together with the difference $\bar{\Phi}_{a}-\Phi_{d}$ (red solid line), as the aggregate result of cars parked in the lot. Again, the proportion of surplus cars is indicated by the dashed-dotted line.
@@
4.3 Optimising Model Parameters from Observed Data